# An Efficient Sign Language Translation Using Spatial Configuration and Motion Dynamics with LLMs

### v1

@@ Abstract
Gloss-free Sign Language Translation (SLT) converts sign videos directly into spoken language sentences without relying on glosses. Recently, Large Language Models (LLMs) have shown remarkable translation performance in gloss-free methods by harnessing their powerful natural language generation capabilities. However, these methods often rely on domain-specific fine-tuning of visual encoders to achieve optimal results. By contrast, this paper emphasizes the importance of capturing the spatial configurations and motion dynamics inherent in sign language. With this in mind, we introduce Spatial and Motion-based Sign Language Translation (SpaMo), a novel LLM-based SLT framework. The core idea of SpaMo is simple yet effective. We first extract spatial and motion features using off-the-shelf visual encoders and then input these features into an LLM with a language prompt. Additionally, we employ a visual-text alignment process as a warm-up before the SLT supervision. Our experiments demonstrate that SpaMo achieves state-of-the-art performance on two popular datasets, PHOENIX14T and How2Sign [^1] Code will be made available upon publication..
An Efficient Sign Language Translation Using Spatial Configuration and Motion Dynamics with LLMs
@@ 1 Introduction
Sign language is a visual means of communication primarily used by Deaf communities, relying on physical movements rather than spoken words. In this paper, we tackle Sign Language Translation (SLT), focusing on converting sign videos into spoken language sentences. Early SLT methods Camgoz et al. (2020); Voskou et al. (2021); Zhou et al. (2021b, a); Yin et al. (2021); Jin et al. (2022); Chen et al. (2022a, b); Zhang et al. (2023b) have primarily relied on glosses—written representations of signs using corresponding words. Glosses provide a structured form of sign language, which helps identify semantic boundaries within continuous sign sequences. This, in turn, allows the models to better comprehend the overall content of the sign videos Yin et al. (2023); Wei and Chen (2023). However, annotating glosses is a labor-intensive and time-consuming process that requires expertise in sign language. This significantly hinders the expansion of sign language datasets and limits the development of SLT methods Li et al. (2020b); Shi et al. (2022); Lin et al. (2023).
@@
To address these limitations, there is a shift towards gloss-free methods that rely solely on the sign videos and corresponding translated text. While these methods still underperform compared to the gloss-based methods, efforts have been made to reduce the performance gap by focusing on temporal semantic structures Li et al. (2020b) and aligning visual and textual modalities Zhao et al. (2021); Yin et al. (2023); Fu et al. (2023); Zhou et al. (2023). Recently, Large Language Models (LLMs) have demonstrated remarkable translation performance in a gloss-free setting by harnessing their powerful language generation capabilities. However, the modality gap between the continuous sign videos and discrete text poses a challenge for the LLMs in effectively understanding the sign videos. To address this problem, these methods fine-tune their visual encoders to be more domain-specific Wong et al. (2024); Chen et al. (2024); Rust et al. (2024); Gong et al. (2024). However, these approaches are resource-intensive and time-consuming, presenting significant challenges that are hard to overcome. Moreover, there has been limited research on how the LLMs process the sign videos and the reasons behind their superior performance compared to other methods.
@@
In this paper, we challenge the understanding that fine-tuning visual encoders is necessary to achieve optimal performance in gloss-free SLT based on LLMs. Instead, we revisit a fundamental question: “What are the key components in sign language that most effectively convey meaning?” Our focus is on the important roles of spatial configuration and motion dynamics in sign language. Spatial configuration encompasses the arrangement and positioning of signs within the signing space, including hand shapes, facial expressions, and body postures. These components work together to distinguish different signs and convey their intended meanings Emmorey and Casey (1995). As shown in Figure 1(a), the signs for “cold” and “winter” both use the same handshape, with a shivering motion of the fists. The primary difference lies in the facial expressions: “cold” is typically accompanied by a tensed or grimaced expression, while “winter” may feature a more neutral expression. Motion dynamics, on the other hand, involve the path, speed, and rhythm of hand movements, illustrating how movements alter the meanings of signs over time Bosworth et al. (2019). As shown in Figure 1(b), the signs for “chair” and “sit” both use the same “H” handshape and involve the interaction of both hands. However, the motion differentiates these signs: “chair” involves a repetitive tapping motion, while “sit” involves a single, smooth motion. These examples highlight the importance of both spatial configuration and motion dynamics in conveying accurate messages in sign language.
@@
To this end, we introduce a novel gloss-free framework, Spatial and Motion-based Sign Language Translation (SpaMo), designed to fully leverage the spatial configurations and motion dynamics in the sign videos, all without the need for domain-specific fine-tuning. As shown in Figure 2, the core idea is simple: We extract spatial and motion features using two different visual encoders and input these features into an LLM with a language prompt. Specifically, we use a pre-trained image encoder (e.g., ViT) as Spatial Encoder (SE), to individually encode each frame for its spatial features. To further refine the spatial configurations, we apply $S^{2}$ scaling Shi et al. (2024), which processes a sign image at multiple scales. Additionally, we use a video encoder (e.g., VideoMAE) as Motion Encoder (ME) to encode sign clips (groups of sign frames) into the motion features. To capture finer motion dynamics, we apply a sliding window approach, which results in implicit gloss-level representations Cheng et al. (2023); Hwang et al. (2024). Next, Sign Adapter (SA), comprising Multi-Layer Perceptron (MLP) layers, transfers these features to the LLM. Additionally, we introduce Visual-Text Alignment (VT-Align), a training strategy to effectively narrow the modality gap, ensuring efficient training and enhanced translation performance.
@@
  • We introduce SPaMo, a novel gloss-free framework. Our method eliminates the need for fine-tuning visual encoders by utilizing readily available, off-the-shelf models. Rather than focusing on the expensive training, it focuses on the fundamental components in sign language, enabling offering a simple yet effective translation with LLMs.
  • Our proposed method achieves state-of-the-art performance on two popular sign language datasets: PHOENIX14T and How2Sign.
  • We provide a novel and comprehensive analysis of how the LLM interprets the sign videos within its embedding space and translate them into corresponding text.
@@ 2.1 Gloss-free Sign Language Translation
Gloss-free SLT directly converts sign videos into spoken language sentences without relying on glosses. These methods, however, often underperform compared to gloss-based methods Camgoz et al. (2020); Voskou et al. (2021); Zhou et al. (2021b, a); Yin et al. (2021); Jin et al. (2022); Chen et al. (2022a, b); Zhang et al. (2023b). To address the performance gap, recent work has focused on several key areas: enhancing the temporal semantic structure Li et al. (2020b), improving the alignment between visual and textual modalities Zhao et al. (2021); Lin et al. (2023); Fu et al. (2023), leveraging large language models (LLMs) Wong et al. (2024); Gong et al. (2024); Chen et al. (2024), and scaling efforts by utilizing larger sign language datasets Uthus et al. (2024); Rust et al. (2024).
@@
Despite these advancements, most gloss-free methods depend on fine-tuning visual encoders using the glosses Li et al. (2020b); Yin et al. (2023); Fu et al. (2023), target translations Zhou et al. (2023); Wong et al. (2024), or self-supervised learning Gong et al. (2024); Rust et al. (2024). In particular, fine-tuning with the glosses (gloss-supervision) helps the visual encoders to be more domain-specific training on continuous or isolated Sign Language Recognition (SLR) datasets, such as WLASL Li et al. (2020a) and PHOENIX14T Camgoz et al. (2018). Consequently, we classify these methods as weakly gloss-free due to the implicit involvement of gloss information, as further elaborated in Section 4.3. On the other hand, the rest of the fine-tuning methods eliminate reliance on these annotations. However, these methods often require substantial resources, making it difficult to achieve robust visual representations and enhance translation performance without access to a sufficiently large dataset. To address this limitation, our approach diverges from this norm by focusing on capturing the spatial configurations and motion dynamics, thereby avoiding the need for resource-intensive fine-tuning.
2.2 Large Language Models
@@
Recently, LLMs Touvron et al. (2023); Chiang et al. (2023); Chung et al. (2024) have demonstrated impressive text generation capabilities, through extensive training on web-scale text corpora. This extensive training has endowed the LLMs with robust generalization abilities across various tasks. Notable applications include multilingual translation Zhu et al. (2023); Zhang et al. (2023a); Gao et al. (2024), pose generation Feng et al. (2024); Zhang et al. (2024a), and, visual question answering Li et al. (2023); Liu et al. (2024a, b). In SLT, LLMs also demonstrate impressive translation performance. These methods focus on aligning high-dimensional visual features with inputs comprehensible to LLMs. To achieve this alignment, the visual encoders are pre-trained to produce language-like tokens Gong et al. (2024), utilize pseudo-glosses Wong et al. (2024), or perform video-grounded text generation tasks Chen et al. (2024).
@@
Our work shows that a domain-specific fine-tuning of the visual encoders is unnecessary for optimal performance. Instead, we extract spatial and motion features and pass them to the LLM through a simple connector, accompanied by a light warm-up process. This approach is both simple and effective, proving that a complex learning process is not required to achieve peak performance.
@@
Figure 2: An overview of the SpaMo framework, which consists of three parts: (i) Sign Feature Extraction: Spatial and motion features are extracted using SE and ME, leveraging the $S^{2}$ and sliding window approaches to capture detailed spatial configurations and motion dynamics. (ii) VT-Align: The extracted features are combined within SA to form a unified sign feature. During training, a warm-up process is employed to ensure that SA has well-initialized weights, effectively bridging the modality gap between the sign video and text. (iii) LLM: Finally, the LLM processes the sign feature along with a language-instructive prompt and is trained using LoRA.
@@ 3 Method
We first give an overview of our framework in Section 3.1, We then explain SE and ME in Sections 3.2 and 3.3, respectively. Next, we discuss SA in Section 3.4 and VT-Align in Section 3.5. Finally, we explain the training details in Section 3.6.
@@ 3.1 Framework Overview
Given a sign video $X=\{x_{i}\}_{i=1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{H\times W}$ represents a frame with height $H$ and width $W$, the objective of SLT is to generate a corresponding spoken language sentence $Y=\{y_{j}\}_{j=1}^{U}$, composed of $U$ words. Previous gloss-free methods Zhou et al. (2023); Wong et al. (2024); Gong et al. (2024); Chen et al. (2024) have involved fine-tuning visual encoders using sign language data to be more domain-specific, thereby improving on translation performance. However, while this fine-tuning process injects more domain knowledge at the feature extraction level, it is often unnecessary and resource-intensive, especially with LLMs, which already maintain rich visual information from the visual encoder in their latent space Zhang et al. (2024b). This creates a trade-off, but we argue that the latter approach is more effective. Thus, we emphasize the importance of encoding the spatial configurations and motion dynamics in sign language and decoding this information through proper alignment and training.
As shown in Figure 2, Spatial Encoder (SE) and Motion Encoders (ME) extract two distinct features from the sign video $X$: Spatial features $Z_{s}$ capture the spatial configurations Emmorey and Casey (1995), and motion features $Z_{m}$ represent the motion dynamics Bosworth et al. (2019). These features are then integrated into a combined sign feature $Z_{sm}$ via Sign Adapter (SA). The combined feature is then fed to an LLM with an language prompt, guiding the LLM to generate the translation in the desired language. Additionally, we perform Visual-Text Alignment (VT-Align) to minimize the gap between the visual and textual modalities before and during training under SLT supervision.
@@ 3.2 Spatial Encoder
SE extracts spatial features $Z_{s}$ from the sign video $X$. We utilize a pre-trained image encoder (e.g., ViT), which is kept frozen, and enhances its capability to capture more detailed spatial information by applying Scaling on Scales ($S^{2}$hi et al. (2024). $S^{2}$ is parameter-free and enables the extraction of multi-scale features without altering the original pre-trained encoder. By processing sign images at multiple resolutions, $S^{2}$ provides a more comprehensive spatial understanding, ensuring that SE captures both fine-grained and broad spatial details for accurate sign language interpretation. The resulting spatial features can be represented as $Z_{s}\in\mathbb{R}^{T\times 2d}$, where $T$ is the number of frames, and $2d$ is the enhanced embedding dimension, reflecting the integration of multi-scale features.
@@ 3.3 Motion Encoder
ME derives motion features from the sign video $X$. Similar to SE, we employ a pre-trained video encoder (e.g., VideoMAE), which remains frozen, to process sign clips segmented from the video. However, accurately segmenting the sign video into distinct gloss-level clips is challenging without the support of pre-trained Continuous Sign Language Recognition (CSLR) models Wei and Chen (2023). To address this limitation, we use a sliding window approach to capture implicit gloss-level representations Cheng et al. (2023); Hwang et al. (2024). Specifically, we divide the sign video into short, overlapping clips, then feed each clip into ME to extract the implicit gloss-level motion features $Z_{m}\in\mathbb{R}^{N\times d}$, where $N=\frac{T}{t}$ and $t$ denotes the number of frames between the start of neighboring clips. Since $Z_{m}$ is generated by processing $N$ short clips, it can also be interpreted as a sequence of $N$ clip-wise features.

### v2

# An Efficient Gloss-Free Sign Language Translation Using Spatial Configurations and Motion Dynamics with LLMs

@@ Abstract
Gloss-free Sign Language Translation (SLT) converts sign videos into spoken language sentences without relying on glosses, which are the written representations of signs. Recently, Large Language Models (LLMs) have shown remarkable translation performance in gloss-free methods by harnessing their powerful natural language generation capabilities. However, these methods often rely on domain-specific fine-tuning of visual encoders to achieve optimal results. By contrast, we emphasize the importance of capturing the spatial configurations and motion dynamics in sign language. With this in mind, we introduce Spatial and Motion-based Sign Language Translation (SpaMo), a novel LLM-based SLT framework. The core idea of SpaMo is simple yet effective: instead of domain-specific tuning, we use off-the-shelf visual encoders to extract spatial and motion features, which are then input into an LLM along with a language prompt. Additionally, we employ a visual-text alignment process as a lightweight warm-up step before applying SLT supervision. Our experiments demonstrate that SpaMo achieves state-of-the-art performance on three popular datasets—PHOENIX14T, CSL-Daily, and How2Sign—without visual fine-tuning  [^1] Code is available at https://github.com/eddie-euijun-hwang/SpaMo.
An Efficient Gloss-Free Sign Language Translation Using Spatial Configurations and Motion Dynamics with LLMs
@@ 1 Introduction
Sign language is a visual means of communication primarily used by Deaf communities, relying on physical movements rather than spoken words. In this paper, we tackle Sign Language Translation (SLT), focusing on converting sign videos into spoken language sentences. Early SLT methods Camgoz et al. (2020); Zhou et al. (2021a); Chen et al. (2022a, b); Zhang et al. (2023b) have primarily relied on glosses—written representations of signs using corresponding words. Glosses provide a structured form of sign language, which helps identify semantic boundaries within continuous sign sequences. This, in turn, allows the models to better comprehend the overall content of the sign videos Yin et al. (2023); Wei and Chen (2023). However, annotating glosses is a labor-intensive and time-consuming process that requires expertise in sign language. This significantly hinders the expansion of sign language datasets and limits the development of SLT methods Li et al. (2020b); Shi et al. (2022); Lin et al. (2023).
@@
To address these limitations, there has been a shift towards gloss-free methods that rely solely on the sign videos and corresponding translations. While these methods still underperform compared to the gloss-based methods, efforts have been made to reduce the performance gap by focusing on temporal semantic structures Li et al. (2020b) and aligning visual and textual modalities Zhao et al. (2021); Yin et al. (2023); Zhou et al. (2023); Zhao et al. (2024). Recently, LLMs have demonstrated remarkable translation performance in a gloss-free setting by harnessing their powerful language generation capabilities. However, the modality gap between the continuous sign videos and discrete text poses a challenge for the LLMs in effectively understanding the sign videos. To address this, many methods fine-tune their visual encoders to be more domain-specific to sign language Wong et al. (2024); Chen et al. (2024); Rust et al. (2024); Gong et al. (2024).
That said, fine-tuning visual encoders can be resource-heavy and time-consuming, making it impractical for real-world applications, especially when considering the diversity of sign languages. This leads to an important question: Is domain-specific tuning of visual encoders necessary to achieve optimal performance in LLM-based SLT? We argue that focusing on the inherent characteristics of sign language could reduce the need for such extensive fine-tuning. First, visual encoders trained on general domains Radford et al. (2021); Oquab et al. (2023) have already proven highly effective in downstream tasks such as action recognition Huang et al. (2024); Tang et al. (2024) and video captioning Yang et al. (2023); Zhou et al. (2024). Moreover, LLMs are capable of retaining rich visual information from these general encoders in their latent space Zhang et al. (2024c). Rather than emphasizing fine-tuning, we shift our attention to the crucial roles of spatial configurations and motion dynamics in sign language.
@@
Spatial configurations include the arranging and positioning of signs within the signing space, including hand shapes, facial expressions, and body postures. These components work together to distinguish different signs and convey their intended meanings Emmorey and Casey (1995). As shown in Figure 1(a), the signs for “cold” and “winter” both use the same handshape, with a shivering motion of the fists. The primary difference lies in the facial expressions: “cold” is typically accompanied by a tense or grimaced expression, while “winter” may feature a more neutral expression. Motion dynamics, on the other hand, involve the path, speed, and rhythm of hand movements, illustrating how movements alter the meanings of signs over time Bosworth et al. (2019). As shown in Figure 1(b), the signs for “chair” and “sit” both use the same “H” handshape and involve the interaction of both hands. However, the motion differentiates these signs: “chair” involves a repetitive tapping motion, while “sit” involves a single, smooth motion. These examples highlight the importance of the spatial configurations and motion dynamics in conveying accurate messages in sign language.
@@
To this end, we introduce a novel gloss-free framework, Spatial and Motion-based Sign Language Translation (SpaMo). SpaMo is designed to fully exploit the spatial configurations and motion dynamics in the sign videos using off-the-shelf visual encoders, all without the need for domain-specific fine-tuning. As shown in Figure 2, the core idea is simple: We extract spatial features (spatial configurations) and motion features (motion dynamics) using two different visual encoders, and feed these into an LLM with a language prompt. Specifically, we use a pre-trained image encoder (e.g., ViT) as Spatial Encoder (SE) to individually encode each frame for its spatial features. To further refine the spatial features, we apply $S^{2}$ scaling Shi et al. (2024), which processes a sign image at multiple scales. Additionally, we use a video encoder (e.g., VideoMAE) as Motion Encoder (ME) to encode sign clips (groups of sign frames) into the motion features. To capture finer motion dynamics, we apply a sliding window approach, which results in implicit gloss-level representations Cheng et al. (2023); Hwang et al. (2024). Next, Sign Adapter (SA), comprising Multi-Layer Perceptron (MLP) layers, transfers these features to the LLM. To further bridge the modality gap, we propose Visual-Text Alignment (VT-Align), a training strategy that aligns the visual features with the LLM’s embedding space, promoting more efficient training and improved translation performance.
@@
  • We introduce SPaMo, a novel gloss-free framework based on LLMs. Our method is simple yet effective, focusing on conveying core elements of sign language to LLMs without domain-specific tuning of visual encoders.
  • Our proposed method achieves state-of-the-art performance on three popular sign language datasets: PHOENIX14T, CSL-Daily, and How2Sign.
  • We provide a novel and comprehensive analysis of how the LLM interprets the sign videos within its embedding space and translates them into corresponding text.
@@ 2.1 Gloss-free Sign Language Translation
Gloss-free SLT directly converts sign videos into spoken language sentences without relying on glosses. These methods, however, often underperform compared to gloss-based methods Camgoz et al. (2020); Zhou et al. (2021b, a); Yin et al. (2021); Chen et al. (2022a, b); Zhang et al. (2023b); Jing et al. (2024). To address the performance gap, recent work has focused on several key areas: enhancing the temporal semantic structure Li et al. (2020b), improving the alignment between visual and textual modalities Zhao et al. (2021); Lin et al. (2023); Fu et al. (2023), leveraging LLMs Wong et al. (2024); Gong et al. (2024); Chen et al. (2024), and scaling efforts by utilizing larger sign language datasets Uthus et al. (2024); Rust et al. (2024). Despite these advancements, most gloss-free methods depend on fine-tuning visual encoders using the glosses Li et al. (2020b); Yin et al. (2023); Fu et al. (2023), target translations Zhou et al. (2023); Wong et al. (2024), or self-supervised learning Gong et al. (2024); Rust et al. (2024). In particular, fine-tuning with the glosses helps the visual encoders to offer more domain-specific training on continuous or isolated Sign Language Recognition (SLR) datasets, such as WLASL Li et al. (2020a) and PHOENIX14T Camgoz et al. (2018).
@@
Consequently, we classify these methods as weakly gloss-free due to the implicit involvement of the glosses, as further elaborated in Section 4.2. On the other hand, the rest of the fine-tuning methods eliminate reliance on these annotations. However, they often require substantial resources. As a results, it can be difficult to achieve robust visual representations and improve translation performance without access to a sufficiently large dataset. To address this limitation, our approach diverges from this norm by focusing on capturing the spatial configurations and motion dynamics through off-the-shelf visual encoders. This allows us to bypass the need for resource-intensive fine-tuning.
2.2 Scaling Language Models in SLT
@@
The scaling laws in language models Kaplan et al. (2020) have been pivotal in the rise of Large Language Models (LLMs) Touvron et al. (2023); Chiang et al. (2023); Chung et al. (2024). Leveraging their strong generation capabilities, LLMs have been applied across diverse domains: multilingual translation Zhu et al. (2023); Zhang et al. (2023a); Gao et al. (2024), pose generation Feng et al. (2024); Zhang et al. (2024b), and visual question answering Li et al. (2023); Liu et al. (2024a, b), extending their impact beyond Natural Language Processing. LLMs have also demonstrated impressive translation performance in the SLT domain. SLT methods using the LLMs focus on aligning high-dimensional visual features with inputs comprehensible to LLMs. These methods involve fine-tuning visual encoders to produce language-like tokens Gong et al. (2024), using pseudo-glosses Wong et al. (2024), or performing video-grounded text generation tasks Chen et al. (2024).
@@
In this work, we take a different approach by focusing on spatial configurations and motion dynamics. We extract spatial and motion features and pass them to LLMs with a light warm-up process. This method is simple yet effective, demonstrating that an extensive pre-training for the visual encoders is unnecessary to achieve peak performance.
@@
Figure 2: An overview of the SpaMo framework, which consists of three parts: (i) Sign Feature Extraction: Spatial and motion features are extracted using SE and ME, using the $S^{2}$ and sliding window approaches to capture detailed spatial configurations and motion dynamics. (ii) VT-Align: The extracted features are combined within SA to form a unified sign feature. During training, a warm-up process is employed to ensure that SA has well-initialized weights, effectively bridging the modality gap between the sign video and text. (iii) LLM: the LLM processes the sign feature along with a language-instructive prompt and is trained using LoRA.
@@ 3 Method
We first give an overview of our framework in Section 3.1. We then explain SE and ME in Sections 3.2 and 3.3, respectively. Next, we discuss SA in Section 3.4 and VT-Align in Section 3.5. Finally, we explain the training details in Section 3.6.
@@ 3.1 Framework Overview
Given a sign video $X=\{x_{i}\}_{i=1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{H\times W}$ represents a frame with height $H$ and width $W$, the objective of SLT is to generate a corresponding spoken language sentence $Y=\{y_{j}\}_{j=1}^{U}$, composed of $U$ words. Previous gloss-free methods Zhou et al. (2023); Wong et al. (2024); Gong et al. (2024); Chen et al. (2024) have involved fine-tuning visual encoders using sign language data to make them more domain-specific, leading to improvements in translation accuracy. However, while this fine-tuning introduces more domain knowledge at the feature extraction level, it is unnecessary, especially with LLMs, which already maintain rich visual information from the visual encoder in their latent space Zhang et al. (2024c). Although there may be a trade-off, we argue that utilizing the spatial configurations and motion dynamics with proper alignment and training, offers a more efficient and effective solution.
As shown in Figure 2, Spatial Encoder (SE) and Motion Encoders (ME) extract two distinct features from the sign video $X$: Spatial features $Z_{s}$ capture the spatial configurations Emmorey and Casey (1995), and motion features $Z_{m}$ represent the motion dynamics Bosworth et al. (2019). These features are then integrated into a combined sign feature $Z_{sm}$ via Sign Adapter (SA). The combined feature is then fed to an LLM with a language prompt, guiding the LLM to generate the translation in the desired language. Additionally, we perform Visual-Text Alignment (VT-Align) to minimize the gap between the visual and textual modalities before and during training under SLT supervision. In the following sections, we provide a detailed explanation of SE, ME, SA, and VT-Align.
@@ 3.2 Spatial Encoder
SE extracts spatial features $Z_{s}$ from the sign video $X$. We utilize a pre-trained image encoder (e.g., ViT), which is kept frozen, and enhances its capability to capture finer spatial information by applying Scaling on Scales ($S^{2}$hi et al. (2024). $S^{2}$ is parameter-free and enables the extraction of multi-scale features without altering the original pre-trained encoder. By processing sign images at multiple resolutions, $S^{2}$ provides a more comprehensive spatial understanding, ensuring that SE captures both fine-grained and broad spatial details for accurate sign language interpretation. The resulting spatial features can be represented as $Z_{s}\in\mathbb{R}^{T\times 2d}$, where $T$ is the number of frames, and $2d$ is the increased embedding dimension, reflecting the integration of multi-scale features.
@@ 3.3 Motion Encoder
ME derives motion features from the sign video $X$. Similar to SE, we employ a pre-trained video encoder (e.g., VideoMAE), which remains frozen, to process sign clips segmented from the video. However, accurately segmenting the sign video into distinct gloss-level clips is challenging without the support of pre-trained Continuous Sign Language Recognition (CSLR) models Wei and Chen (2023). To address this limitation, we use a sliding window approach to capture implicit gloss-level representations Cheng et al. (2023); Hwang et al. (2024). Specifically, we divide the sign video into short, overlapping clips, then feed each clip into ME to extract the implicit gloss-level motion features $Z_{m}\in\mathbb{R}^{N\times d}$, where $N$ is the number of segments. The number of segments $N$ is calculated as $N=\left\lfloor\frac{T-w}{s}\right\rfloor+1$, where $T$ is the total number of frames, and $w$ and $s$ are the window size and stride, respectively. Since $Z_{m}$ is generated by processing $N$ short clips, it can also be interpreted as a sequence of $N$ clip-wise features.